# How In-Context Learning Emerges from Training on Unstructured Data: On the Role of Co-Occurrence, Positional Information, and Noise Structures

### v1

@@ Appendix A Proof of Theorem 1
Let $|V|=2K+L$ denote the vocabulary size. Consider a sentence $X$ represented by its one-hot encoding (i.e., $X\in\{0,1\}^{|V|\times S}$). For every position $i\in[S]$, the loss for predicting the word in the $i$-th position given all the other words is given by $||AX(\mathds{1}_{S}-e_{i})-Xe_{i}||_{2}^{2},$ where $A=\frac{U^{\top}V}{S-1}\in\mathbb{R}^{|V|\times|V|}$ and $e_{i}\in\mathbb{R}^{S}$ is a zero vector with $1$ on its $i$-th entry. Here, $A$ is a matrix summarizing the similarity between each pair of words, one as a center word and the other as a context word. Our objective is to find $A$ that minimizes the sum of losses for each position in each sentence. Lemma 8 gives a closed-form expression of the minimizer.
@@
completing the proof. ∎
Appendix B Proof of Theorem 2
@@
The next step the proof is to use Lemma 8 in Appendix A to obtain the similarity matrix $A$. As previously, we denote a square matrix with $\alpha$ on the diagonal and $\beta$ off the diagonal as $X_{\alpha,\beta}$, and a matrix with all entries $\gamma$ as $Y_{\gamma}$. We then have
@@
for some $q_{1},q_{2},\cdots,q_{10}$. Recall that our task is show that given context words $c_{i_{1}},d_{i_{1}},\cdots,c_{i_{\ell}},d_{i_{\ell}}c_{i_{\ell+1}}$ with distinct $i_{k}$’s, the center word is more likely to be $d_{i_{\ell+1}}$ than $e_{i_{\ell+1}}$. In other words, we need to establish that
@@
Appendix C Proof of Theorem 3
@@
Appendix D Proof of Proposition 4
@@
Appendix E Proof of Theorem 5
@@
  • When the first five tokens do not contain any noise token, output a uniform probability vector over $\mathcal{N}$.
  • Otherwise, output the conditional probability of $c[2]$ given $x$, where $(x,c[2])\in S_{1}$. Here, $x$ represents the last non-noise token.
@@
  • When the seventh token is not a noise token, output the seventh token with probability one.
  • When the seventh token is a noise token, output a uniform probability vector over $\mathcal{N}$.
Under this rule, the predicted output for any in-context example abacdcef is e, resulting in perfect ICL accuracy. ∎
Appendix F Proof of Theorem 6
@@
Appendix G Proof of Theorem 7
@@
Appendix H Details of experiments and data sets
@@
All experiments employ the Adam optimizer [Kingma and Ba, 2015] with a learning rate of $0.001$. Early stopping is applied based on validation loss with a patience threshold of 5, utilizing a randomly selected subset representing 50% of the original data set. All transformer models use two heads, as increasing the number of heads does not appear to significantly impact performance.
@@
The world_population.csv data set, used for the experiments in Sections 2.4 and 2.5, is obtained from Kaggle. According to the author, this data set is created from World Population Review.
@@
The us-state-capitals.csv data set, used for the experiments in Section 2.4, is obtained from this Github repository. Its source is unclear.
@@
The uscities.csv data set, used for the experiments in Section 2.4, is obtained from Simple Maps, with a CC 4.0 license.

### v2

@@ Appendix B Proof of Theorem 1
Let $|V|=2K+L$ denote the vocabulary size. Consider a sentence $X$ represented by its one-hot encoding (i.e., $X\in\{0,1\}^{|V|\times S}$). For every position $i\in[S]$, the loss for predicting the word in the $i$-th position given all the other words is given by $||AX(\mathds{1}_{S}-e_{i})-Xe_{i}||_{2}^{2},$ where $A=\frac{U^{\top}V}{S-1}\in\mathbb{R}^{|V|\times|V|}$ and $e_{i}\in\mathbb{R}^{S}$ is a zero vector with $1$ on its $i$-th entry. Here, $U$ ($V$) is a matrix consisting of the center (context) embeddings of all tokens, and $A$ is a matrix summarizing the similarity between each pair of words (one as a center word and the other as a context word). Our objective is to find $A$ that minimizes the sum of losses for each position in each sentence. Lemma 8 gives a closed-form expression of the minimizer.
@@
completing the proof.
Note that this condition ensures that the model predicts $d_{\ell+1}$ instead of one of the $r_{i}$’s. ∎
Appendix C Comparison of ICL performance using squared and cross-entropy loss across different numbers of examples
Table 8: ICL performance in the clean scenario, evaluated with both squared and cross-entropy loss functions across different numbers of examples (0 to 8) with $d_{E}=100$, averaged over 10 repetitions.
Squared Cross-entropy
$(p_{0},p_{1},p_{2})$ 0 2 4 6 8 0 2 4 6 8
$(0,1,0)$ 1 1 0 0 0 0.87 0 0 0 0
$(0,0,1)$ 1 1 1 0 0 1 0 0 0 0
$(1/2,1/2,0)$ 1 1 1 1 1 1 1 0.34 0 0
$(1/2,0,1/2)$ 1 1 1 1 1 1 1 1 1 1
$(0,1/2,1/2)$ 1 1 1 1 1 1 1 1 0 0
$(1/3,1/3,1/3)$ 1 1 1 1 1 1 1 1 1 0
Table 9: ICL performance in the corrupted scenario, evaluated with both squared and cross-entropy loss functions across different numbers of examples (0 to 8) with $d_{E}=100$, averaged over 10 repetitions.
Squared Cross-entropy
$(p_{0},p_{1},p_{2})$ 0 2 4 6 8 0 2 4 6 8
$(0,1,0)$ 1 0 0 0 0 0 0 0 0 0
$(0,0,1)$ 1 0.97 0 0 0 1 0 0 0 0
$(1/2,1/2,0)$ 1 1 1 0.53 0 1 0 0 0 0
$(1/2,0,1/2)$ 1 1 1 1 1 1 1 1 1 1
$(0,1/2,1/2)$ 1 1 0.76 0 0 1 1 0 0 0
$(1/3,1/3,1/3)$ 1 1 1 1 1 1 1 1 0.18 0
From Tables 8 and 9, we observe that ICL with CBOW on single-relationship tasks performs better with squared loss compared to cross-entropy loss and with fewer demonstration examples. Also, ICL tends to deteriorate after a certain number of in-context demonstrations. As detailed in Appendix B, a smaller number of examples (e.g., zero) allows the model to produce the correct output instead of one of the $r_{i}$’s. This is in contrast with transformer-based LLMs, which achieve better ICL performance as the number of demonstrations increases. On the other hand, ICL on dual-relationship tasks as described in Section 2.2 requires at least one demonstration example to distinguish between the two tasks.
Appendix D Proof of Theorem 2
@@
The next step the proof is to use Lemma 8 in Appendix B to obtain the similarity matrix $A$. As previously, we denote a square matrix with $\alpha$ on the diagonal and $\beta$ off the diagonal as $X_{\alpha,\beta}$, and a matrix with all entries $\gamma$ as $Y_{\gamma}$. We then have
@@
for some $q_{1},q_{2},\cdots,q_{10}$. Recall that we want to show that given $c_{i_{1}},d_{i_{1}},\cdots,c_{i_{\ell}},d_{i_{\ell}}c_{i_{\ell+1}}$ with distinct $i_{k}$’s as context words, the center word is more likely to be $d_{i_{\ell+1}}$ than $e_{i_{\ell+1}}$. In other words, we need to establish that
@@
Appendix E Proof of Theorem 3
@@
Appendix F Proof of Proposition 4
@@
Appendix G Proof of Theorem 5
@@
  • When the first five tokens do not contain any nuisance token, output a uniform probability vector over $\mathcal{N}$.
  • Otherwise, output the conditional probability of $c[2]$ given $x$, where $(x,c[2])\in S_{1}$. Here, $x$ represents the last non-nuisance token.
@@
  • When the seventh token is not a nuisance token, output the seventh token with probability one.
  • When the seventh token is a nuisance token, output a uniform probability vector over $\mathcal{N}$.
Under this rule, the predicted output for any in-context example abacdcef is e, resulting in perfect ICL accuracy.
Appendix H Proof of Theorem 6
@@
Appendix I Proof of Theorem 7
@@
Appendix J Details of experiments and data sets
J.1 Architecture and implementation
@@
All experiments utilize the Keras package in Python, employing the Adam optimizer [29] with a learning rate of 0.01. Early stopping is applied based on validation loss with a patience threshold of 5, utilizing a randomly selected subset representing 50% of the original data set. Each transformer layer uses two heads, as we empirically demonstrated that increasing the number of heads does not impact performance in our experiments. Each layer consists of the following components (in order): (1) Keras’ multi-head causal self-attention block, with key_dim $=$ value_dim $=$ embed_dim$/2$; (2) Skip connection and layer normalization; (3) One hidden layer feed-forward network using the ReLU activation with dimension =2×=2$\times$= 2 × embed_dim; and (4) Skip connection and layer normalization.
J.2 Source and details of data sets
@@
The world_population.csv data set, used for the experiments in Section 2, is obtained from Kaggle. According to the author, this data set is created from World Population Review.
@@
The us-state-capitals.csv data set, used for the experiments in the beginning of Section 2, is obtained from this Github repository. Its source is unclear.
@@
The uscities.csv data set, used for the experiments in the beginning of Section 2, is obtained from Simple Maps, with a CC 4.0 license.
J.3 Details of synthetic data used in experiments
Below we provide additional details regarding the synthetic data used in our experiments.
  1. 1.
    For experiments in Table 1, the training data consists of 50,000 sentences. In the clean version, sentences are generated uniformly as described in Section 2.1. In the corrupted version, sentences are generated in a similar manner, but each $(c_{i},d_{i})$ pair is replaced by $(c_{i},r_{j})$ or $(d_{i},r_{j})$ with a probability of $1/4$ each. Test sentences are generated according to the setup in Theorem 1. Some examples are as follows:
    • Clean
      • Training: $c_{1}d_{1}r_{1}r_{2}r_{3}r_{4}r_{5}r_{6}$ or $r_{1}r_{2}r_{3}r_{4}r_{5}r_{6}r_{7}r_{8}$
      • Prompt: $c_{1}d_{1}c_{2}d_{2}c_{3}d_{3}c_{4}?$
    • Corrupted
      • Training: $c_{1}r_{1}r_{2}r_{3}r_{4}r_{5}r_{6}r_{7}$ or $c_{1}d_{1}c_{2}r_{1}r_{2}r_{3}r_{4}r_{5}$
      • Prompt: $c_{1}d_{1}c_{2}d_{2}c_{3}d_{3}c_{4}?$
  2. 2.
    For experiments in Table 2, the training data consists of 50,000 sentences. In the clean version, sentences are generated uniformly as described in Section 2.2. In the imbalanced and extreme versions, the 60 other words are divided into three categories: 20 for $cd$ sentences ($rcd_{\cdot}$), 20 for $ce$ sentences ($rce_{\cdot}$), and 20 for both types ($r_{\cdot}$). In the imbalanced version, $cd$ ($ce$) sentences are 4 times more likely to sample a $cd$ ($ce$) word than a $ce$ ($cd$) word. In the extreme version, $cd$ ($ce$) sentences cannot contain any $ce$ ($cd$) words. Test sentences are generated according to the setup in Theorem 2. Some examples are as follows:
    • Clean examples
      • Training: $c_{1}d_{1}r_{1}r_{2}r_{3}r_{4}r_{5}r_{6}$ or $c_{1}e_{1}r_{1}r_{2}r_{3}r_{4}r_{5}r_{6}$
      • Prompt: $c_{1}d_{1}c_{2}d_{2}c_{3}d_{3}c_{4}?$ or $c_{1}e_{1}c_{2}e_{2}c_{3}e_{3}c_{4}?$
    • Imbalance examples
      • Training: $c_{1}d_{1}rcd_{1}rcd_{2}rcd_{3}rce_{4}r_{5}r_{6}$ or $c_{1}e_{1}rcd_{1}r_{2}rce_{3}rce_{4}rce_{5}r_{6}$
      • Prompt: $c_{1}d_{1}c_{2}d_{2}c_{3}d_{3}c_{4}?$ or $c_{1}e_{1}c_{2}e_{2}c_{3}e_{3}c_{4}?$
    • Extreme examples
      • Training: $c_{1}d_{1}rcd_{1}rcd_{2}rcd_{3}r_{4}r_{5}r_{6}$ or $c_{1}e_{1}r_{1}r_{2}r_{3}rce_{4}rce_{5}r_{6}$
      • Prompt: $c_{1}d_{1}c_{2}d_{2}c_{3}d_{3}c_{4}?$ or $c_{1}e_{1}c_{2}e_{2}c_{3}e_{3}c_{4}?$
  3. 3.
    Experiments in Table 3 follow the setup of experiments in Table 2, except that the pairs are now of the form $(c_{i},d_{i})$ and $(e_{i},f_{i})$ instead of $(c_{i},d_{i})$ and $(c_{i},e_{i})$.
  4. 4.
    For experiments in Section 2.4, the corpus generation process is as follows:
      Example: Gabon (GAB) protects its diverse rainforests and wildlife.
    • Generate 60 sentences without any country, capital city, or IOC code.
    The ICL prompts follow the form used in the country-capital city and US state-capital city experiments in the beginning of Section 2, with 1 to 5 in-context examples.
  5. 5.
    For experiments in Table 4, the training and test data consist of all sentences in the form $abca$, where $a$, $b$, and $c$ are distinct. Each test sentence is different from any training sentence. In the first scenario (both), the first tokens of the training sentences cover the entire vocabulary. In the second scenario (either), each token can be the first token in either the training or test data, but not both.
  6. 6.
    For experiments in Table 5, the training data consists of 50,000 sentences generated uniformly as detailed in Section 3.1. The ICL prompt formats are also described in Section 3.1.
  7. 7.
    For experiments in Table 6, the training data consists of 50,000 sentences. In the clean scenario, the training data are of the form $abcadefd$ and $abcbdefe$, with ICL prompts as $\overline{abcadef}?$ and $\overline{abcbdef}?$. In the block-noisy scenario, the training data include sequences like $n_{1}n_{2}n_{3}n_{4}abcadefd$ and $abcbn_{1}n_{2}n_{3}n_{4}defe$, with ICL prompts as $\overline{abcadefdghi}?$ and $\overline{abcbdefeghi}?$.
  8. 8.
    For experiments in Table 7, the training data consists of 50,000 sentences generated uniformly according to the processes in Sections 4.1 and 4.2. The ICL prompt formats are also described in the same subsections.
  9. 9.
    For experiments in Section 4.3, the corpus generation process is as follows:
    • Generate 130 sentences containing exactly one country-capital pair (13 for each country).
      Example: Paramaribo stands as capital of Suriname.
    • Generate 60 sentences without any country, capital city, or IOC code.